# Glonet: Mercator’s end-to-end neural forecasting system

### v1

@@ 1 Introduction
The accurate prediction of oceanic states is crucial for numerous marine operations, encompassing navigation [James, 1957], fisheries management [Pikitch et al., 2004], disaster response [Teal and Howarth, 1984, Breivik et al., 2013, Zohdi and Abbaspour, 2019], and climate research. Ocean forecasting has primarily depended on physics-based numerical ocean models [Rosati and Miyakoda, 1988, Madec et al., 1997, Chassignet et al., 2007] co-developed and operated by various operational forecasting centers worldwide. These numerical models employ intricate systems of partial differential equations to simulate ocean dynamics, sometimes with the constraint of not departing too far from both contemporary and historical oceanographic observations. Operational Global Ocean Forecasting Systems (GOFSs) merge these numerical frameworks with real-time observational database via advanced data assimilation techniques [Bannister, 2017, Brasseur et al., 2005, Cummings and Smedstad, 2013, Lellouche et al., 2018, Copernicus, 2023] to deliver precise and timely forecasts. However, these models face significant challenges: they are computationally expensive, slow to improve, and require substantial resources for forecasting the state of the ocean, limiting their ability to provide forecasts rapidly, especially in situations that require high-frequency updates or providing ensemble forecasts and their associated uncertainty. This sluggishness also hampers their use in research, where exploring a wide range of scenarios quickly and efficiently can be essential for advancing ocean science. Moreover, even with the influx of more observational data, it is often challenging to improve these models due to their reliance on fixed physical assumptions and the need for extensive parameter tuning. The need for faster, more flexible approaches becomes then evident as traditional models struggle to adapt to evolving requirements and increasingly diverse data sources [Liu et al., 2017].
In recent years, the weather forecasting community has witnessed a significant surge in the development and deployment of neural network (NN)-based forecasting systems. These data-driven models leverage advancements in machine learning, particularly deep learning, to enhance forecast accuracy and computational efficiency. Weather forecasting models such as FourCastNet [Pathak et al., 2022], GraphCast [Lam et al., 2022], Pangu-Weather [Bi et al., 2023] and recently AIFS [Lang et al., 2024] have demonstrated remarkable improvements in prediction performance and speed, often outperforming traditional numerical weather prediction (NWP) methods in various metrics [Rasp et al., 2024]. The rapid adoption of NN-based systems in meteorology underscores the potential of artificial intelligence (AI) to revolutionize environmental forecasting.
@@
Addressing these challenges, Mercator Ocean International (MOI) has developed a novel neural network-based forecasting system, GLONET, which represents a significant advancement in the field of ocean forecasting. GLONET is trained on the $1/12^{\circ}$ daily mean outputs of the global Mercator Ocean physical reanalysis GLORYS12 [Lellouche et al., 2021] and is designed to operate at a horizontal resolution of $1/4^{\circ}$ while inheriting fine-scale dynamics of GLORYS12, enabling it to resolve mesoscale oceanic features with unprecedented accuracy. Leveraging a hierarchical transformer-based backbone, GLONET incorporates physics-based principles through neural operators and networks to dynamically capture local-global interactions within a unified, scalable framework, mitigating the complexities introduced by continental boundaries, and ensuring high small-scale accuracy and efficient dynamics over short forecast intervals.
To ensure the operational viability of GLONET, MOI has established robust pre-operational pipelines that facilitate its integration into existing forecasting workflows. Additionally, GLONET undergoes rigorous validation using the Class-4 framework [Ryan et al., 2015, Hernandez et al., 2009, Divakaran et al., 2015, Lellouche et al., 2013], a stringent evaluation protocol that assesses forecast accuracy against leading operational GOFSs. In addition to traditional metrics, GLONET also undergoes a unique NN-specific evaluation criterion, proposed in this paper, to meet the high standards required for operational deployment. In addition, using the aforementioned validation frameworks, GLONET is benchmarked against the current global Mercator Ocean analysis and forecasting $1/12^{\circ}$ high-resolution physical system GLO12 [Lellouche et al., 2023], and Xihe, one of the first neural global ocean forecasting system [Wang et al., 2024] trained from the same ocean reanalysis GLORYS12.
@@ 2 Model
A diverse range of deep learning architectures such as vision transformers, graph neural networks, and neural operators has found application in data-driven weather forecasting and, more recently, in ocean forecasting [Wang et al., 2024], where vision and Swin transformers [Khan et al., 2022, Liu et al., 2021] were combined to develop a global ocean forecasting system. Building on these advances, we introduce GLONET, a multi-scale forecasting system. The ocean is fundamentally a turbulent system, marked by the interplay of processes across a wide range of spatial and temporal scales. To address this complexity, GLONET leverages Fourier neural operators [Li et al., 2020] to model broad, basin-wide patterns like gyres, currents, tele-connection and global circulation, capturing large-scale dependencies effectively representing the large-scale energy transfer that drives the background state of the ocean. At the small-scale dynamics, GLONET employs convolutional neural networks (CNNs) [O’Shea, 2015] to enhance predictions of submesoscale features, which influence localized mixing and fine-scale dependencies. Additionally, GLONET employs a encoder-decoder [Schmidhuber, 2015] to fuse these multi-scale circulations into a unified latent space, capturing essential detail across different scales that play a key role in redistributing heat, salt, and energy.
@@
where $\mu$ is a probability measure over the initial conditions function space $\mathcal{X}$.
@@
where $\mathcal{C}_{\psi}:\mathcal{F}\rightarrow\mathcal{U}$ is a convolutional neural networks parameterized by $\psi$, and $\mathcal{U}$ the feature space capturing localized dependencies. This part enhances the model’s ability to predict detailed, high-resolution oceanic phenomena that are critical for accurate localized forecasts. To integrate the large-scale and small-scale feature representations, $\mathcal{G}_{\theta}$ employs an encoder-decoder that fuses multi-scale circulations from $\mathbf{v}_{t+1}$ and $\mathbf{u}_{t+1}$ into a unified latent representation:
@@
3 Training
GLONET is trained on the daily mean outputs of the GLORYS12 reanalysis and is configured to produce daily forecasts by using the ocean state inputs as shown in Figure 2. To optimize computational efficiency, GLONET operates on a coarser $[1/4]^{\circ}$ grid, by interpolating GLORYS12 products to this resolution. Despite this reduction in grid resolution, the model is designed to retain and represent the small-scale ocean dynamics intrinsic to the original $[1/12]^{\circ}$ products. This approach allows GLONET to balance the computational tractability of daily forecasts with the need to accurately capture mesoscale and finer processes, thus providing a nuanced representation of ocean dynamics without the high computational cost associated with finer grids.
Variable Depth Level [m]
Temperature 0.49, 47, 92, 155,
Salinity 222, 318, 380, 453, 541, in/out
U and V current components 643, 763, 902, 1245, 1684
2225, 3220, 3597, 3992
4405, 4833, 5274 $m$
SSH Surface in/out
Figure 2: GLONET’s configuration, characterized by $[1/4]^{\circ}$ of spatial resolution and 21 vertical levels ranging from $0.4$ to $5274$ $m$ of depth. U and V correspond to the zonal and meridional velocity respectively. SSH corresponds to the sea surface height.
@@ 3 Training
Training was performed on $1/4^{\circ}$ interpolated GLORYS12 daily mean outputs for the period $[1993,2019]$, using Adam’s optimizer [Kingma, 2014], and a learning rate decreasing from $e^{-4}$ to $e^{-5}$. Once the model is trained and the desired performances were reached, a second round of fine-tuning is performed to optimize the forecasts accurary up to 4 days. The overall training took around 3 weeks while using a total of 32 40G-A100 GPUs. GLONET is capable of outputting 10 days forecast in less than 10 seconds.
@@ 4 Results
In this section, GLONET’s forecasting capabilities are empirically assessed using a comprehensive set of validation metrics, designed to capture different dimensions of forecast quality. The evaluation begins with the IV-TT Class 4 Framework, a robust benchmark that leverages observational data to evaluate forecasting accuracy against reference datasets. This is followed by point-wise accuracy assessments, process-oriented metrics, and diagnostic analyses, offering a multifaceted understanding of the model’s performance. Point-wise evaluations use GLORYS12 as the reference dataset, assessing the models’ ability to generate forecasts that align closely with the data on which they were trained. This analysis provides a detailed view of the models’ accuracy in capturing ocean states across various spatial and temporal scales, ensuring consistency and comparability with the training reference. Extending beyond these localized evaluations, process-oriented metrics are incorporated to evaluate the model’s ability to derive essential ocean quantities and maintain coherence across interconnected forecasted variables. Additionally, diagnostic variables, such as vorticity and mesoscale eddies are analysed to assess the model’s skill in preserving dynamic consistency, energy cascading and fine-scale structures. This holistic evaluation ensures that GLONET’s forecasts not only minimize numerical errors but also adhere to the fundamental dynamics governing ocean systems.
@@ 4.1 IV-TT Class 4 Framework Evaluation
In this part, GLONET, Xihe and GLO12 are evaluated using the CLASS4 dataset, developed and maintained by the "OceanPredict Task Team for Intercomparison and Validation" (IV-TT team). The CLASS4 framework provides a benchmark for model validation by operating within the observation space, enabling a direct comparison between observed and modeled values across both spatial and temporal dimensions. For each observation, the corresponding model counterpart is extracted at the same spatial and temporal location across various forecast lead times, ranging from the best analysis (day 0) to ten-day forecasts. The CLASS4 dataset includes observations of temperature and salinity from Argo profiles, sea surface temperature (SST) from surface drifting buoys, and sea level anomaly (SLA) from along-track satellite measurements. This framework serves as a robust tool for intercomparison, facilitating a comprehensive assessment of the forecasting models’ performance.
Figure 3: Dispersion and evolution of RMSD as a function of forecast lead time for temperature (top left) and salinity (top right) in the 5–100 m layer, SLA (bottom left), and SST (bottom right) from drifting buoys. The thick lines represent the $75\%$ distribution, while the thin lines correspond to the $95\%$ distribution.
Figure 3 illustrates the dispersion and evolution of the Root Mean Square Difference (RMSD) as a function of forecast lead time for the following variables: temperature and salinity (5–100 m layer), SLA, and SST. XIHE performs comparably to GLO12 and better than GLONET in forecasting temperature in the 5–100 m layer. The three models exhibit similar performance in salinity, though GLONET shows greater dispersion for the 3- and 5-day forecasts. GLONET surpasses XIHE in SLA, consistently outperforming GLO12 for forecasts ranging from 5 to 9 days. For SST, GLO12 demonstrates superior performance across all forecast horizons, except for the one-day forecast, where the three models yield identical results.
@@
(a) GLO12 vs GLONET
(b) GLO12 vs XIHE
Figure 4: Normalized MAE difference between GLO12 and GLONET (left) and between GLO12 and XIHE (right). Positive values (red) denote improvement relative to GLO12, while negative values (blue) indicate degradation.
The neural network-based models, GLONET and XIHE, generally perform well in dynamically active regions such as the Kuroshio, Gulf Stream, Antarctic Circumpolar Current (ACC), and parts of the North Atlantic and North Pacific. However, both models show reduced performance in the tropical band and within subtropical gyres. GLONET outperforms XIHE in the South Pacific, potentially due to better generalization in regions with sparse observations or lower variability in training data. Conversely, XIHE demonstrates slightly better performance in the North Atlantic, likely benefiting from denser observational data in this well-monitored region.
@@ 4.2 Point-wise evaluation
In this part, the point-wise forecasting accuracy of GLONET is evaluated by comparing its outputs against GLORYS12 reanalysis products, which serve as the reference. GLORYS12 provides high-quality, observationally informed ocean state estimates, making it an established benchmark for assessing forecast performance in ocean modeling. To contextualize GLONET’s forecasting capabilities within the domain of data-driven modeling, its 10-day forecasts are compared with those of Xihe, another NN-based forecasting system. Both GLONET and Xihe are evaluated using identical initial conditions, the same GLORYS12 reference fields, and over a consistent evaluation period spanning January to July 2024. This setup ensures a fair and uniform comparison, enabling an objective assessment of GLONET’s performance relative to other data-driven approaches.
Figure 5: Root Mean Square Error (RMSE) averaged across all vertical levels for 3D variables (U and V currents, temperature, salinity), alongside the RMSE of sea surface height (SSH), for GLONET (orange line) and Xihe (blue line). The calculations cover the period from January to July 2024, using GLORYS12 as the reference dataset. Forecasts are initialized weekly on Wednesdays based on nowcast analyses performed with GLO12, lagging real-time by seven days.
Figure 5 displays the forecasting performance of GLONET and Xihe, expressed in terms of Root Mean Square Error (RMSE). The RMSE is calculated for all forecasted variables over a 10-day forecast window, averaged across all vertical levels for 3D variables, for the period from January to July 2024. In general, GLONET produces highly skillful forecasts as can be seen from the RMSE scores, particularly for ocean currents compared to Xihe, with an improvement of $23\%$.
A gradual decrease of forecast accuracy over time is observed for GLONET, mainly due to its auto-regressive fashion. Conversely, Xihe exhibits relatively stable performance across all lead times, which can be attributed to its design, where separate models are trained for each specific lead time.
The RMSE scores presented here represent averages across all vertical levels for 3D variables. Since the models operate with distinct vertical level structures, the depth-wise accuracy of forecasts is also investigated. This time, all 10-day forecasted lead times are averaged for the period of January to July 2024, for each vertical level. This allows us to examine the depth-wise accuracy of the forecasts for a more granular understanding of the different models’ performances.
@@
Figure 6: RMSE computed at each depth level and averaged over all lead times for 3D variables (U and V currents, temperature, and salinity) for the GLONET (orange line) and Xihe (blue line) models, covering January to July 2024. GLORYS12 serves as the reference, with 10-day forecasts initialized weekly on Wednesdays from a nowcast analysis performed with GLO12 seven days behind the real-time.
@@
Figure 6 shows the RMSE scores averaged over all lead times for each vertical level, highlighting variations in forecasting accuracy at different depths. This analysis reveals that, in terms of ocean currents, GLONET continues to perform well, with Xihe coming closer to GLONET’s performance in the upper levels, particularly within the mixed layer. For temperature, GLONET performs comparably to Xihe in the intermediate layers. However, GLONET diverges slightly from Xihe at the surface and bottom levels. This depth-dependent analysis provides valuable insights into the models’ forecasting capabilities at different ocean vertical levels, illustrating their strengths and weaknesses across the vertical profile.
@@
In addition to the above analyses, the 10-day forecast accuracy is further evaluated by averaging RMSE across only the common depth levels shared by all models as shown in Figure 7. This analysis reveals that GLONET continues to perform well in predicting ocean currents, closely matching Xihe’s accuracy within the mixed layer. For temperature, GLONET’s performance is similar to Xihe at intermediate depths, while Xihe shows slight advantages at the surface and the bottom layer.
Figure 8: Spatial distribution of RMSE at the surface level for GLONET and Xihe across all forecasted variables (U and V currents, temperature, and salinity and SSH). The RMSE is computed with respect to GLORYS12 over the evaluation period from January to July 2024. These maps provide insight into the regional performance of each model, highlighting areas of stronger or weaker agreement with the reference reanalysis. Results are averaged over all lead times for weekly forecasts initialized on Wednesdays, adhering to the operational protocol.

### v2

# GLONET: Mercator’s end-to-end neural Global Ocean forecasting system

@@ Abstract
[…]
Machine Learning and Computation
Mercator Ocean International, Toulouse, France.
[…]
Anass El Aouniaelaouni@mercator-ocean.eu
{keypoints}
GLONET combines physics-based principles with neural networks to effectively capture local-global ocean interactions.
A series of comprehensive validation metrics is proposed, specifically tailored for neural network-based ocean forecasting systems.
GLONET’s experimental daily forecasts are accessible through the European Digital Twin Ocean platform EDITO.
Plain Language Summary
Accurate ocean forecasting is vital for various applications, from scientific research to decision-making in marine operations. Traditional forecasting systems rely on complex physics-driven numerical models, which can be computationally expensive and slow to improve. GLONET offers a data-driven alternative by utilizing machine learning to predict ocean conditions efficiently. By learning from historical ocean data, GLONET captures intricate ocean dynamics, providing accurate and timely forecasts. Its performance has been validated against existing forecasting systems, showing promising results in predicting key ocean variables.
@@ 1 Introduction
[…]
@@
[…]
@@ 2 Model
[…]
To formalize the learning process, we frame GLONET within an operator learning perspective, which naturally suits the spatio-temporal nature of ocean forecasting. Instead of directly predicting the next ocean state through fixed-size feature transformations, we view forecasting as learning a mapping between function spaces: the space of initial conditions to the space of future states. This formulation allows GLONET to inherently respect the continuous, spatially varying structure of ocean fields, and aligns with the recent trend of modeling physical systems as operators acting on entire fields rather than isolated grid points. This perspective is crucial for modeling the ocean’s interconnected behavior, where changes at one location can influence distant regions through wave propagation, advection, and teleconnections.
@@
where $\mu$ is a probability measure over the initial conditions function space $\mathcal{X}$. In practical terms, this theoretical formulation guides the design of GLONET’s architecture by emphasizing the need for global context, multi-scale feature extraction, and flexible handling of spatially distributed inputs. Rather than treating prediction as a local, pixel-wise task, the operator-based view motivates the explicit modeling of broad spatial dependencies (via Fourier neural operators) alongside fine-grained local dynamics (via CNNs). This dual approach is crucial for improving forecast accuracy in complex systems like the ocean, where processes operate and interact across multiple scales simultaneously.
@@
where $\mathcal{C}_{\psi}:\mathcal{F}\rightarrow\mathcal{U}$ is a convolutional neural networks parameterized by $\psi$, and $\mathcal{U}$ the feature space capturing localized dependencies. The combination of Fourier-based global modeling and CNN-based local refinement directly addresses the dual nature of ocean dynamics: the need to accurately forecast both broad circulation patterns and localized phenomena such as eddies and fronts. Without the operator-based view, such explicit separation and integration of scales would be much harder to achieve within a single model.
To integrate the large-scale and small-scale feature representations, $\mathcal{G}_{\theta}$ employs an encoder-decoder that fuses multi-scale circulations from $\mathbf{v}_{t+1}$ and $\mathbf{u}_{t+1}$ into a unified latent representation:
@@
3 Data and Training
3.1 Data
[…]
3.2 Training
GLONET is trained on daily mean outputs from the GLORYS12 reanalysis and is configured to produce daily forecasts using the ocean state variables listed in Table 1. The selected vertical levels are designed to sample key dynamical regimes across the full ocean column. The first four levels, located at approximately 0.49 m, 47 m, 92 m, and 155 m, reside within the upper 200 meters and capture near-surface variability, the seasonal mixed layer, and the upper thermocline, regions central to air-sea interactions, mixed layer dynamics, and stratification onset. The intermediate levels (222 m to 541 m) sample the deeper thermocline and upper intermediate waters, where vertical gradients remain pronounced and contribute to processes such as subsurface ventilation and water mass transformation. The subsequent range from 643 m to 1245 m spans the lower thermocline into the upper deep ocean, where vertical variability decreases but key features, including the base of the thermocline and intermediate water masses, continue to exert dynamical influence. Levels deeper than 1245 m, extending to 5274 m, are included to represent the structure of the abyssal ocean and to account for large-scale circulation patterns, stratification, and bathymetric effects.
While the direct impact of deep ocean conditions on upper-ocean forecasts at a 10-day timescale is expected to be limited, the inclusion of full-depth profiles during training allows the model to learn cross-depth statistical relationships and to better capture low-frequency variability such as seasonal cycles and basin-scale dynamics. Additionally, by exposing the model to deep ocean structure, it is indirectly informed about the influence of bathymetry and topographic constraints on the general circulation, particularly in regions where bottom features steer currents or shape mesoscale variability. This vertically comprehensive configuration supports better generalization and enhances the physical consistency of the forecasts within a unified, data-driven framework.
To optimize computational efficiency, GLONET operates on a coarser $[1/4]^{\circ}$ grid, by interpolating GLORYS12 products to this resolution. This decision is motivated by the substantial increase in memory requirements associated with high-resolution spatiotemporal models, where the memory footprint grows approximately quadratically with the spatial resolution. Moreover, during training, additional memory overhead arises from the need to store intermediate activations for backpropagation and gradient computation, further exacerbating resource demands. Interpolating the inputs to $[1/4]^{\circ}$ thus enables efficient training and inference while preserving the key mesoscale dynamics intrinsic to the original $[1/12]^{\circ}$ products, ensuring a balance between computational tractability and spatial fidelity.
Importantly, in physics-based numerical models, the grid resolution imposes a fundamental constraint on the smallest scales that can be dynamically resolved, with coarse resolutions inherently limiting the representation of mesoscale and submesoscale processes. In contrast, machine learning models trained on interpolated high-resolution data are not governed by the same dynamical restrictions. In the present configuration, interpolating the $[1/12]^{\circ}$ GLORYS12 fields to a $[1/4]^{\circ}$ grid preserves a substantial portion of the small-scale variability, as confirmed through both spectral analyses and qualitative assessments. This stands in contrast to physical models natively run at $[1/4]^{\circ}$ resolution, where the absence of finer grid spacing results in a significant smoothing of the fields and a degradation of mesoscale features. Consequently, the interpolation strategy employed in GLONET enables the model to remain exposed to rich mesoscale dynamics during training, despite operating at a coarser computational resolution, thereby maintaining a critical balance between computational efficiency and physical realism.
Variable Depth Levels [m] Spatial Resolution
Temperature 0.49, 47, 92, 155 $[1/4]^{\circ}$ (input/output)
Salinity 222, 318, 380, 453, 541 $[1/4]^{\circ}$ (input/output)
U and V current components 643, 763, 902, 1245, 1684, $[1/4]^{\circ}$ (input/output)
2225, 3220, 3597, 3992, 4405, 4833, 5274
SSH Surface $[1/4]^{\circ}$ (input/output)
Table 1: GLONET configuration: input/output variables, their corresponding depth levels, and spatial resolution. U and V correspond to the zonal and meridional velocity components, respectively. SSH corresponds to the sea surface height.
@@ 3.2 Training
The training set consists of daily mean GLORYS12 reanalysis outputs from 1993 to 2019. The year 2020 is reserved for validation, providing an unseen dataset for tuning hyperparameters and monitoring generalization during training. For the final evaluation, GLONET is tested using independent initial conditions derived from the operational forecast products GLO12 to rigorously assess out-of-sample performance. During training, we employ random shuffling of samples to enhance statistical robustness and prevent potential temporal or spatial biases within each batch, ensuring more efficient gradient updates and faster convergence.
@@
[…]
@@ 4 Results
In this section, GLONET’s forecasting capabilities are empirically assessed using a comprehensive set of validation metrics, designed to capture different dimensions of forecast quality. The evaluation begins with the IV-TT Class 4 Framework, a robust benchmark that leverages observational data to evaluate forecasting accuracy against reference datasets. This is followed by point-wise accuracy assessments, process-oriented metrics, and diagnostic analyses, offering a multifaceted understanding of the model’s performance.
The models used in this evaluation differ significantly in their design and application. GLONET, a data-driven model, employs an autoregressive structure, generating forecasts based on prior predictions. In contrast, Xihe, also a data-driven approach, does not rely on autoregressive modeling and instead focuses on independent, single-step forecasts, which provides stable performance across lead times but may lack long-term consistency. GLO12, a physics-driven model, integrates governing physical equations and observational data assimilation, allowing it to perform well in dynamically consistent predictions. Each model differs in resolution and computational efficiency, with GLONET and Xihe benefiting from a more flexible, data-driven architectures that can be trained to resolve high-dimensional inputs, while GLO12 rely on more traditional, computationally intensive physics-based and autoregressive fashion.
Model Type Autoregressive Computational Efficiency Resolution
GLONET Data-driven Yes few seconds (GPU) $1/4^{\circ}$
Xihe Data-driven No few seconds (GPU) $1/12^{\circ}$
GLO12 Physics-based Yes 1h (HPC) $1/12^{\circ}$
Table 2: Comparison of model types, autoregressive features, computational efficiency, and resolution.
Point-wise evaluations use GLORYS12 as the reference dataset, assessing the models’ ability to generate forecasts that align closely with the data on which they were trained. This analysis provides a detailed view of the models’ accuracy in capturing ocean states across various spatial and temporal scales, ensuring consistency and comparability with the training reference. Extending beyond these localized evaluations, process-oriented metrics are incorporated to evaluate the model’s ability to derive essential ocean quantities and maintain coherence across interconnected forecasted variables. Additionally, diagnostic variables, such as vorticity and mesoscale eddies are analysed to assess the model’s skill in preserving dynamic consistency, energy cascading and fine-scale structures. This holistic evaluation ensures that GLONET’s forecasts not only minimize numerical errors but also adhere to the fundamental dynamics governing ocean systems.
@@ 4.1 IV-TT Class 4 Framework Evaluation
In this part, GLONET, Xihe and GLO12 are evaluated using the CLASS4 dataset, developed and maintained by the ”OceanPredict Task Team for Intercomparison and Validation” (IV-TT team). The CLASS4 framework provides a benchmark for model validation by operating within the observation space, enabling a direct comparison between observed and modeled values across both spatial and temporal dimensions. For each observation, the corresponding model counterpart is extracted at the same spatial and temporal location across various forecast lead times, ranging from the best analysis (day 0) to ten-day forecasts.
The CLASS4 dataset includes observations of temperature and salinity from Argo profiles, sea surface temperature (SST) from surface drifting buoys, sea level anomaly (SLA) from along-track satellite measurements and surface current observations at 15m depth from the Global Drifter Program (GDP) drifters buoys. This framework serves as a robust tool for intercomparison, facilitating a comprehensive assessment of the forecasting models’ performance. For SST, it is important to note that comparisons are made between the observed SST (from surface drifting buoys) and the modeled SST at the first vertical level (which is at 0.49m depth) for each model. This level is the closest available approximation of the sea surface in the model outputs.
The evaluation of model performance is based on two common metrics: the Root Mean Square Difference (RMSD) and the Mean Absolute Error (MAE).
The equation for RMSE is given by:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$ (10)
where $y_{i}$ represents the observed value, $\hat{y}_{i}$ the modeled value, and $N$ is the total number of data points.
Similarly, the MAE is calculated as:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|$ (11)
where $|y_{i}-\hat{y}_{i}|$ is the absolute difference between the observed and modeled values.
Figure 2 illustrates the dispersion and evolution of the Root Mean Square Difference (RMSD) as a function of forecast lead time for the following variables: temperature, salinity (5–100 m layer), SLA, SST, zonal and meridional surface currents. XIHE performs comparably to GLO12 and slightly better than GLONET in forecasting temperature of the 5–100 m layer. The three models exhibit similar performance in salinity, though GLONET shows greater dispersion for the 3 day and 5 day forecasts. GLONET surpasses XIHE in SLA and surface currents, consistently outperforming GLO12 for forecasts ranging from 5 to 9 days. For SST, GLO12 demonstrates superior performance across all forecast horizons, except for the one-day forecast, where the three models yield identical results. For surface currents, GLONET achieves the highest performance in both meridional and zonal components, indicating that global ocean circulation is well captured in its surface layers relative to GDP observations at 15 m depth.
Figure 2: Dispersion and evolution of RMSD as a function of forecast lead time for salinity (top left) and temperature (top middle) in the 5–100 m layer, SLA (top right), and SST/(Temperature of the first level for models) (bottom left) from drifting buoys and zonal and meridional ocean currents (middle of bottom right). The thick lines represent the $75\%$ distribution, while the thin lines correspond to the $95\%$ distribution, the dot represent the median of the distribution
@@
Figure 3: Normalized MAE difference for SST/(Temperature of the first level for models) and zonal current between GLO12 and GLONET (left panels) and XIHE (righ panels). Positive values (red) denote improvement relative to GLO12, while negative values (blue) indicate degradation.
For SST the neural network-based models, GLONET and XIHE, generally perform well in dynamically active regions such as the Kuroshio, Gulf Stream, Antarctic Circumpolar Current (ACC), and parts of the North Atlantic and North Pacific. However, both models show reduced performance in the tropical band and within subtropical gyres. GLONET outperforms XIHE in the South Pacific, potentially due to better generalization in regions with sparse observations or lower variability in training data. Conversely, XIHE demonstrates slightly better performance in the North Atlantic, likely benefiting from denser observational data in this well-monitored region.
For surface currents, GLONET demonstrates robust performance across most regions, with mild localized performance reductions observed in the North and South Pacific Gyre and equatorial zones. GLONET exhibits superior accuracy compared to XIHE in all major ocean basins, including the Pacific, Atlantic, and Indian Oceans. Notably achieving exceptional results in dynamically active regions, such as the Antarctic Circumpolar Current (ACC), where it captures strong current features with high fidelity.
@@ 4.2 Point-wise evaluation
In this part, the point-wise forecasting accuracy of GLONET is evaluated by comparing its outputs against GLORYS12 reanalysis products, which serve as the reference. GLORYS12 provides high-quality, observationally informed ocean state estimates, making it an established benchmark for assessing forecast performance in ocean modeling.
To contextualize GLONET’s forecasting capabilities within the domain of data-driven modeling, its 10-day forecasts are compared with those of Xihe, another neural network-based forecasting system. Both GLONET and Xihe are evaluated using identical initial conditions, the same GLORYS12 reference fields, and over a consistent evaluation period spanning January to July 2024. This setup ensures a fair and uniform comparison, enabling an objective assessment of GLONET’s performance relative to other data-driven approaches.
The evaluation of point-wise accuracy is based on the Root Mean Square Error (RMSE), a standard metric for quantifying the difference between observed and predicted values. The RMSE is calculated for each forecast lead time and for each of the model variables (temperature, salinity, and surface current components). For each model and each variable, the RMSE is computed as:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$ (12)
where $y_{i}$ are the observed values from the GLORYS12 dataset, $\hat{y}_{i}$ are the forecasted values from the model (GLONET or Xihe), and $N$ is the number of data points over the evaluation period.
This RMSE-based evaluation allows for a precise comparison of the forecasting accuracy of GLONET and Xihe, highlighting differences in their ability to replicate observed ocean states across various temporal scales and forecast lead times.
@@
Figure 4: RMSE computed at each depth level and averaged over all lead times for 3D variables (U and V currents, temperature, and salinity) for the GLONET (orange line) and Xihe (blue line) models, covering January to July 2024. GLORYS12 serves as the reference, with 10-day forecasts initialized weekly on Wednesdays from a nowcast analysis performed with GLO12 seven days behind real-time.
@@
Figure 4 shows the RMSE scores averaged over all lead times at each vertical level, highlighting variations in forecasting accuracy with depth. In terms of ocean currents, GLONET consistently outperforms Xihe across most depths, while at surface and mixed-layer depths, the performances are comparable. For temperature, GLONET and Xihe exhibit similar accuracies in the intermediate ocean layers, with slight differences at the surface and bottom levels. This depth-resolved evaluation provides a more detailed understanding of each model’s strengths and weaknesses across the vertical profile.
@@
To ensure a fairer comparison given differences in vertical discretizations between models, RMSE scores are also averaged across the common depth levels shared by GLONET and Xihe, as shown in Figure 5. This analysis confirms that GLONET maintains strong performance for ocean currents throughout the forecast window. In terms of temperature, GLONET and Xihe perform similarly at intermediate depths, while Xihe shows a slight advantage near the surface and at the bottom layers.
Figure 6: Spatial distribution of RMSE at the surface level for GLONET (left) and Xihe (right) across all forecasted variables (U and V currents, temperature, salinity, and SSH). The RMSE is computed with respect to GLORYS12 over the evaluation period from January to July 2024. Results are averaged over all lead times for weekly forecasts initialized on Wednesdays, adhering to the operational protocol.